6. Consider a sample X1, ... ,X, U(0,0) and answer following questions. (a) Suggest a reasonable estimator of 8 and explain. hi i am struggling with this problem, thanks
Answer #1

ABW) X1, X2, ---, Xn Lady uro, o) 8 (4): 3: 4 likelihood function of sample values, La frui) = (0) We arange sample observat(b) Y= max(* 182, - - ,Xn) codifi of Y = Fyly) = Pe Cysy) = Pe[max. of (X1 X2, -- , Xn) sy] = Pe[ x5 y, x = y, ----,Xnsy] = P20(49)-(130) Saha ep met onze - Recet; 4. ) =613 comment of the) - (R)3 I- TAMI Itu 4 - Itu u hp-taff ch! =(hja 05h50 methu =Var (4) = E(yº) – (E(y)]2 - 69.) 02 (67) 0] Var (Y) = 02 [(#2) - (n. 1)] = Answer (d) y = max (X1, X2, -- , Xn) smice Ely) =is an unbiased estimator thus ô = (mt) y Var (O) = var [(ht!) y] = (hu? you (4) var () = (*#1 )? [CM+2) - 67.)? 3.02 standard